### v1

# SDTrack: A Baseline for Event-based Tracking via Spiking Neural Networks

Abstract
Event cameras provide superior temporal resolution, dynamic range, power efficiency, and pixel bandwidth. Spiking Neural Networks (SNNs) naturally complement event data through discrete spike signals, making them ideal for event-based tracking. However, current approaches that combine Artificial Neural Networks (ANNs) and SNNs, along with suboptimal architectures, compromise energy efficiency and limit tracking performance. To address these limitations, we propose the first Transformer-based spike-driven tracking pipeline. Our Global Trajectory Prompt (GTP) method effectively captures global trajectory information and aggregates it with event streams into event images to enhance spatiotemporal representation. We then introduce SDTrack, a Transformer-based spike-driven tracker comprising a Spiking MetaFormer backbone and a simple tracking head that directly predicts normalized coordinates using spike signals. The framework is end-to-end, does not require data augmentation or post-processing. Extensive experiments demonstrate that SDTrack achieves state-of-the-art performance while maintaining the lowest parameter count and energy consumption across multiple event-based tracking benchmarks, establishing a solid baseline for future research in the field of neuromorphic vision.
Figure 1: Comparison of SDTrack with other trackers on VisEvent in terms of energy consumption, AUC and model size.
1 Introduction
Visual object tracking is a fundamental task in computer vision, with extensive applications in surveillance, autonomous driving, and robots. Conventional RGB-based tracking methods have seen considerable advancements [4, 6, 5, 35, 40, 47], but often fail under extreme conditions. In contrast, event cameras present attractive advantages, including higher temporal resolution (in the order of $\mu s$) and higher dynamic range (140dB vs. 60dB). These features enable event cameras to locate objects in degraded conditions. Recently, Spiking Neural Networks (SNNs) have gathered increasing attention due to their sparse and spike-driven computation paradigm. Compared with traditional Artificial Neural Networks (ANNs), spiking neurals transmit information through sparse binary spikes, i.e., 0 or 1, which enable them to replace computationally intensive multiply-accumulate (MAC) in ANNs with efficient energy-efficient accumulation (AC) operations. This makes SNNs naturally align with the sparse feature of event data and the low power consumption properties of event cameras. Consequently, the integration of SNNs into event-based tracking emerges as a compelling and synergistic approach.
Several studies have applied SNNs to event-based tracking with promising results. However, most of these works adopt a hybrid architecture that combines ANNs and SNNs, and fail to implement cross-correlation operations between template and search features, which has been proven to enhance performance. Consequently, current SNN-based trackers neither fully exploit the energy-efficiency advantages inherent to SNNs nor achieve their maximum performance capabilities. To address these challenges, this paper presents the first transformer-based Spike-Driven tracking (SDTrack) pipeline, which achieves state-of-the-art (SOTA) performance in the field of event-based tracking through fully spike-driven operations. The proposed pipeline consists of an innovative event data processing method capable of extracting trajectory information and a transformer-based fully spike-driven tracker.
For the SDTrack pipeline’s input, we observe that existing event aggregation methods exhibit insufficient capacity to generate robust representations for real-time event flow. To address this limitation, we propose a Global Trajectory Prompt (GTP) method that effectively captures trajectory information embedded within the event flow. Specifically, GTP accumulates positive and negative polarities in the first two channels respectively, while recording trajectory information in the third channel. This three-channel image representation of event data not only provides enhanced temporal information representation but also effectively aligns with data formats commonly used in computer vision tasks, facilitating better utilization of pre-trained weights for training the tracker.
For the fully spike-driven tracker in our pipeline, we implement a three-aspect design, including position information acquisition, backbone, and tracking head. Firstly, we design a novel Intrinsic Position Learning (IPL) method, which endows the network with the ability to autonomously learn positional information, unlike existing trackers that require explicit position encoding steps. Secondly, we introduce the MetaFormer architecture [41, 42] to event-based tracking for the first time, constructing a streamlined spike-driven backbone. This backbone employs depth-wise separable convolutions for the convolutional components, while utilizing only vanilla attention in the transformer modules. Thirdly, we provide a fully spike-driven tracking head. Notably, our tracker operates in an end-to-end manner, without relying on any data augmentation or complex post-processing operations.
We conduct extensive experiments to access our methods. On one hand, we validate the efficacy and generality of our GTP method, providing new baselines for mainstream trackers across multiple event-based benchmarks. On the other hand, we demonstrate that SDTrack achieves an excellent balance among parameter count, energy consumption, and accuracy, as depicted in Fig.1. The main contributions of this paper are summarized as follows:
  • We propose a novel event aggregation method called GTP, which effectively captures global trajectory information, yielding enhanced spatiotemporal representation of event flow. This method enables diverse trackers to achieve substantial performance gains.
  • We first propose a transformer-based spike-driven tracker called SDTrack, which consists of a novel IPL method, a streamlined backbone, and a spike-driven tracking head. SDTrack operates in an end-to-end manner, without relying on complex pre- and post-processing operations.
  • Extensive experiments show that SDTrack achieves SOTA results in multiple benchmarks with reduced parameter and energy consumption. These results demonstrate the potential of SNNs in event-based tracking tasks and establish a baseline for future research.
2 Related Works
Event Aggregation Method in Tracking. To leverage advanced trackers designed for traditional cameras, researchers typically divide asynchronous event flow into sub-event streams of equal length and aggregate them into synchronous event images [44, 30, 28, 43, 29, 45, 18, 24, 49, 31]. However, current event aggregation methods in the tracking domain are suboptimal. The Event Frame method only records the latest polarity at each spatial location [44, 28], losing valuable information when multiple changes occur at a pixel. While Time-Surface [18] and Event Count [24] methods can implicitly represent multi-directional motion, their temporal representation lacks robust trajectory information for tracking. Zhu et al.[49] and Wang et al.[31] aggregate latest occurrence times of polarity changes, but their four-channel representation performs poorly in complex tasks and limits transfer learning, as pre-trained weights are designed for three-channel inputs [14, 13]. Therefore, we propose GTP, which captures both short-term multi-directional movements and global trajectory information, enhancing spatiotemporal representation while maximizing transfer learning potential.
Figure 2: Overview of SDTrack Pipeline. After receiving both template and search event streams, the GTP processes each into event frames, which are then concatenated into a unified matrix by IPL. Following SNN Conv Block processing, the matrix undergoes restoration and tokenization. The template features, after cross-correlation computation with search features, are fed into the SNN Tracking Head for target position and scale prediction. The detailed design of each module is illustrated in the middle and the bottom of the figure.
Event-based Tracking. Various studies [29, 28] demonstrate that trackers in RGB-based single object tracking (SOT) perform inadequately on event-based tracking tasks. Research indicates this deficiency stems from the lack of texture information needed by RGB-based trackers in the event data [44]. Given the sparse spike-driven nature of SNNs, several studies suggest there exists an inherent compatibility between SNNs and event data. Consequently, several studies has employed SNNs to construct event-based trackers [27, 1, 11] or tracking systems for edge devices [22, 16, 36]. Among these SNN-based trackers, STNet [44] and SNNTrack [46] that utilize SNNs to extract temporal features have emerged as competitive trackers. However, both of them implement hybrid architectures combining SNNs and ANNs, and neither leverages self-attention for template-search interaction. This results in the low-power benefit and performance potential of SNNs still not being fully utilized. To address these limitations, we design the first transformer-based spike-driven tracker, aimed at leveraging the inherent energy efficiency advantage of SNNs while fully unleashing their performance potential.
3 SDTrack Pipeline
In this section, we first present the spiking neuron, which constitutes the basic unit of the SDTrack pipeline. Second, we propose an event aggregation method with enhanced representation capabilities, termed GTP. Third, we introduce a transformer-based spike-driven tracker, comprised of the intrinsic position learning, backbone, and head.
3.1 I-LIF Spiking Neuron
Recently, the emergence of I-LIF neurons enhances the expressiveness of SNNs [23, 34, 20, 25, 39]. This enhancement enables SNNs to potentially outperform ANNs in various machie learning tasks [23, 20, 34], while preserving their spike-driven characteristics. Therefore, we employ the I-LIF spiking neuron model to build our SDTrack pipeline. I-LIF neurons transmit information through spike discharge frequency in the training stage, while utilizing binary spike signals during the inference stage [23, 39]. In the training stage, the output of the I-LIF neuron can be expressed as
${\mathbf{s}^{\ell}}=\frac{1}{T}\cdot\left\lfloor\operatorname{clip}\left\{{\mathbf{x}^{\ell}},\text{0},T\right\}\right\rceil,$ (1)
[…]
In the inference stage, the firing rate $\mathbf{s}^{\ell}$ is decoupled into $T$ binary spike signals, mathematically described as
${\mathbf{s}^{\ell}}=\frac{1}{T}\cdot\sum\limits_{t=1}^{T}{\mathbf{s}^{\ell}[t]}.$ (2)
At this stage, for neurons in layer $l+1$, whether considering matrix multiplication or Hadamard product as linear operators, their input can be expressed as
${\mathbf{x}^{\ell+1}}={\mathbf{w}^{\ell}}\cdot{\mathbf{s}^{\ell}}=\sum\limits_{t=1}^{T}{\frac{{{\mathbf{w}^{\ell}}}}{T}}{\mathbf{s}^{\ell}}[t],$ (3)
where $\mathbf{w}^{\ell}$ is the weight matrix in linear or convolutional layer, ${\mathbf{s}^{\ell}}[t]\in\{0,1\}$ is a binary spike at time $t$. Therefore, I-LIF maintains the key spike-driven property of SNNs during inference, which enables the transformation of MAC into efficient AC operations, significantly reducing storage and computational overhead during edge deployment.
Figure 3: Example of event stream and event aggregation. The upper shows an event stream of a suspended object moving rightward then leftward from center in a short time. The lower displays results using various event aggregation methods for this stream.
3.2 Global Trajectory Prompt
Event cameras respond to brightness changes asynchronously and independently for each pixel. For each pixel, when the change in log-scale intensity exceeds a threshold, it outputs an event $e=\{x,y,t,p\}$, where $x$ and $y$ are spatial coordinates, $t$ is the timestamp, and $p$ is the polarity. To leverage rich research achievements in the vision model, various studies aggregates event flow within a time interval into event images that can be processed by vision models. The proposed GTP adopts the same idea, with the difference being that the event images aggregated by GTP have three features: (1) multi-directional motion capture, (2) global trajectory information, and (3) three-channel representation. Specifically, for the $i$-th event image, GTP computes the pixel-wise accumulation in the first two channels using the following equations
$h_{i}^{1}(x,y)\buildrel\textstyle.\over{=}\alpha\cdot\sum\limits_{{t_{k}}\in L}{\delta(x-{x_{k}},y-{y_{k}})\delta({p_{k}}-1)},$ (4)
$h_{i}^{2}(x,y)\buildrel\textstyle.\over{=}\alpha\cdot\sum\limits_{{t_{k}}\in L}{\delta(x-{x_{k}},y-{y_{k}})\delta({p_{k}}+1)},$ (5)
where $L$ is the duration of each event stream and $\alpha$ is a coefficient that preserves more valid data while removing partial noise. By using our GTP, the first and second channels in a event image, i.e., $h_{i}^{1}$ and $h_{i}^{2}$, accumulate the number of positive and negative polarities within the event stream, respectively. As shown in Fig 3, when multiple directional movements occur at a pixel within the interval $L$, prior methods only select the most recent polarity value as the aggregation result, discarding preceding motion information. In contrast, GTP separately accumulates the number of positive and negative polarities, thus preserving all motion information throughout the entire time interval.
In addition to recording all motion information via the first two channels, GTP captures global trajectory information in the third channel of the event image, described as
$h_{i}^{3}(x,y)\buildrel\textstyle.\over{=}h_{i-1}^{3}(x,y)\cdot\beta+\alpha\cdot\sum\limits_{j=1}^{2}{C(h_{i-1}^{j}(x,y),h_{i}^{j}(x,y))},$ (6)
$C(h_{i-1}^{j},h_{i}^{j})\buildrel\textstyle.\over{=}\delta(h_{i-1}^{j}(x,y)-0)\cdot[1-\delta(h_{i}^{j}(x,y)-0)],$ (7)
where $\delta$ denotes the Kronecker delta function, and $\beta$ is a decay factor. The $h_{0}^{3}$ is set to a zero matrix. As described above, when starting to accumulate the third channel of the i-th event image, we decay each pixel in the previous event image $h_{i-1}^{3}$ by a factor $\beta$, serving as initial information for the current image’s third channel. The aggregation result for the third channel is obtained upon completion of the entire event stream interval. As depicted in Fig. 3, GTP method to successfully capture the global trajectory information embedded within event data, thereby providing cues for object position and shape determination for tracking.
Overall, the $i$-th event image is constructed by stacking $h_{i}^{1}$, $h_{i}^{2}$, and $h_{i}^{3}$ to form a three-channel representation, which aligns with the RGB format prevalent in conventional vision tasks. As a result, our GTP method enables the accumulated event image to maintain compatibility with existing computer vision architectures, thereby facilitating the utilization of pre-trained weights for tracking applications.
3.3 Transformer-based Spike-driven Tracker
Our SDTrack appears in the yellow section at the top of Fig. 2. Note our processing of its inputs. This subsection describes the design details of SDTrack.
Intrinsic Position Learning. Motivated by two key observations. First, joint positional encoding of template and search images enhances their cross-correlation performance [17]. Second, residual convolutional blocks preceding SNN-based Transformer architectures enable effective learning of positional information [48, 38, 37]. Based on these insights, we propose Intrinsic Position Learning (IPL). In SDTrack pipeline, IPL concatenate the event-aggregated template image $\mathbf{Z}\in(T,C,{H_{z}},{W_{z}})$ and search image $\mathbf{X}\in(T,C,{H_{x}},{W_{x}})$ diagonally before feeding them into the SNN Conv Module
$\mathbf{U}=\operatorname{IPL}(\mathbf{X},\mathbf{Z}),$ (8)
where $\mathbf{U}\in(T,C,{H_{z}+H_{x}},{W_{z}+W_{x}})$. As demonstrated in Tab. 2, our approach enables SDTrack to capture stronger positional information compared to conventional encoding schemes, without explicitly requiring positional encoding. Due to the spike-driven nature of our architecture, the additional padding introduces negligible computational overhead.
Backbone. The backbone of SDTrack consists of a SNN Conv Module containing several SNN Conv Blocks and a SNN Transformer Module containing several SNN Transformer Blocks. Specifically, the SNN Conv Block can be written as
$\quad\mathbf{U}^{\prime}=\mathbf{U}+\operatorname{SNNSepConv}(\mathbf{U})$ (9)
$\quad\mathbf{U}^{\prime\prime}=\mathbf{U}^{\prime}+\operatorname{SNNConvGroup}(\mathbf{U}^{\prime})$ (10)
$\quad\operatorname{SNNConvGroup}(\mathbf{U}^{\prime})=f(f(\mathbf{U}^{\prime})),$ (11)
[…]
$\quad\mathbf{V}_{1}=\operatorname{Conv_{pw}}(\mathcal{SN}(\mathbf{U})),$ (12)
$\quad\mathbf{V}_{2}=\operatorname{Conv_{dw}}(\mathcal{SN}(\mathbf{V}_{1})),$ (13)
$\quad\operatorname{SNNSepConv}(\mathbf{U})=\operatorname{Conv_{pw}}(\mathcal{SN}(\mathbf{V}_{2})),$ (14)
where $\mathcal{SN}(\cdot)$ is the I-LIF spiking neuron layer, $\operatorname{Conv_{pw}}(\cdot)$ and $\operatorname{Conv_{dw}}(\cdot)$ are pointwise convilution and depthwise convolution[7]. $f$ denotes the $\operatorname{Conv(\mathcal{SN}(\cdot))}$ operation. Note, we ignore the batch normalization layer for ease of writing.
Before feeding them into the SNN Transformer Module, they are split back to their original positions prior to entering the SNN Conv Module and tokenized. SNN Transformer Block can be written as
$\mathbf{U}^{\prime}=\mathbf{U}+\operatorname{SSA}(\mathbf{U}),$ (15)
$\mathbf{U}^{\prime\prime}=\mathbf{U}^{\prime}+\operatorname{SNNMLP}(\mathbf{U}^{\prime}).$ (16)
The SNN Self-Attention (SSA) module enables tokens with high-level features from both template and search regions to interact through correlation operations, extracting target-oriented features. The input token sequence is transformed into Query ($\mathbf{Q_{s}}$), Key ($\mathbf{K_{s}}$), and Value ($\mathbf{V_{s}}$) spikes through three learnable linear matrices. The correlation operation process can be described as
$\operatorname{SSA}(\mathbf{Q_{s}},\mathbf{K_{s}},\mathbf{V_{s}})=\mathbf{Q_{s}}{\mathbf{K_{s}}^{\top}}\mathbf{V_{s}}*s,$ (17)
where $s$ is a scaling factor that is set in relation to the input channel dimensions and the number of attention heads.
Tracking Head. Similar to many trackers, SDTrack uses a simple center tracking head to directly predict the target position and size. What distinguishes SDTrack is its SNN-redesigned center head. In each branch, the first four convolutional operations are based on spike signals while the final convolutional operation is based on floating-point signals. This design addresses the inherent limitation of spike-based convolutions, which are insufficient for precise object localization and scale estimation—a deficiency that cannot be compensated through continuous learning. We demonstrate the necessity of this approach for constructing SNN-based trackers through both experimental validation in Sec 5 and theoretical derivation in the Supplementary Materials. Furthermore, in Sec 4.4, we demonstrate that center heads are more suitable for SDTrack compared to corner heads.
Training and Inference. Similar to prior works, we train SDTrack through image pair matching [40, 4, 35]. The loss computation incorporates both classification and regression losses. We adopt the weighted focal loss [19] for classification. With the predicted bounding box, $\mathcal{L}1$ loss and the generalized IoU loss [26] are employed for bounding box regression. The overall loss function is
$\mathcal{L}={\mathcal{L}_{\operatorname{cls}}}+{\lambda_{\operatorname{iou}}}{\mathcal{L}_{\operatorname{iou}}}+{\lambda_{\mathcal{L}1}}{\mathcal{L}_{1}},$ (18)
where ${\lambda_{iou}}=2$ and ${\lambda_{L1}}=5$ are the regularization parameters in our experiments, following [35]. The detailed description of the loss function is provided in the Supplementary Material. Since neurons transmit spike firing frequencies during the training stage, the network’s simulation time step is set to 1.
During inference, we follow the standard SOT procedure by using the first frame of the test sequence as the template. Notably, we do not employ dynamic template strategies or post-processing techniques such as hanning window penalty. Our GTP method is applicable to real-world scenarios in real-time, as illustrated by the purple box in Fig. 2. Notably, during inference, the network’s simulation time step is set to T, as the spike firing rate used in training is systematically decoupled into precise spike signals over T time steps according to Eq. 2.
Methods Parameters (M) Timestep Power (mJ) FE108 [43] FELT [29] VisEvent [28]
AUC(%) PR(%) AUC(%) PR(%) AUC(%) PR(%)
STARK [35] 28.23 1 58.88 57.4 89.2 39.6 51.7 34.1 46.8
SimTrack [4] 88.64 1 93.84 56.7 88.3 36.8 47.0 34.6 47.6
OSTrack256 [40] 92.52 1 98.90 54.6 87.1 35.9 45.5 32.7 46.4
ARTrack256 [32] 202.56 1 174.80 56.6 88.5 39.5 49.4 33.0 43.8
SeqTrack-B256 [6] 90.60 1 302.68 53.5 85.5 33.0 42.0 28.6 43.3
HiT-B [17] 42.22 1 19.78 55.9 88.5 38.5 48.9 34.6 47.6
GRM [12] 99.83 1 142.14 56.8 89.3 37.2 47.4 33.4 47.7
HIPTrack [3] 120.41 1 307.74 50.8 81.0 38.2 48.9 32.1 45.2
ODTrack [47] 92.83 1 335.80 43.2 69.7 29.7 35.9 24.7 34.7
SiamRPN* [21] - 1 - - - - - 24.7 38.4
ATOM* [9] - 1 - - - 22.3 28.4 28.6 47.4
DiMP* [2] - 1 - - - 37.8 48.5 31.5 44.2
PrDiMP* [10] - 1 - - - 34.9 44.5 32.2 46.9
MixFormer* [8] 37.55 1 - - - 38.9 50.4 - -
STNet* [44] 20.55 3 - - - - - 35.0 50.3
SNNTrack* [46] 31.40 5 8.25 - - - - 35.4 50.4
SDTrack-Tiny 19.61 4 8.16 59.0 91.3 39.6 50.1 35.6 49.2
SDTrack-Base 107.26 4 30.52 59.9 91.5 39.8 50.7 37.4 51.5
Table 1: Comparison with standard SOT pipeline on three event-based tracking benchmarks. Due to incompatibility with the GTP methodology, ODTrack utilizes the data processing approach proposed by Wang et al. for comparative analysis [28]. Works marked with * indicate direct adoption of results from their respective benchmark reports. The remaining results employ the GTP methodology for event aggregation to better demonstrate SDTrack’s performance advantages. The energy consumption of SDTrack is represented by its measurements on the VisEvent dataset. The best two results are shown in red and blue fonts.
4 Experiments
4.1 Implement Details
[…]
We first pretrain the backbone of SDTrack on the ImageNet-1K dataset. We then fine-tune it on the corresponding event-based tracking datasets through a pair matching task [35, 4, 40, 17]. During this process, no data augmentation or preprocessing is utilized. The hyperparameters for the fine-tuning process depend on the dataset and model size. Complete details of pretraining and fine-tuning are available in the Supplementary Materials. All experiments are conducted on 4 * NVIDIA RTX 4090 GPUs. We will open-source all available models and training code.
4.2 Evaluation Metric
We conduct comparative analyses with existing methods primarily across four metrics: Area Under the Curve (AUC), Precision Rate (PR), parameter count, and theoretical energy consumption. The AUC is computed from the area under the Success Plot curve, while PR represents the percentage of samples where the distance between the predicted and ground truth target centers is below 20 pixels.
For direct encoding [33] and to achieve stronger representational capacity for position information and target size, the first layer of SDTrack and the final layer of each group convolutional of tracking head inevitably utilize floating-point-based convolution. All other computations are spike-driven, enabling the substitution of MAC operations with AC operations. Therefore, the theoretical energy consumption of SDTrack can be calculated as
$E$ $=T\cdot({E_{\operatorname{MAC}}}\cdot\sum\limits_{i}{\operatorname{FL}_{\operatorname{SNNConv}}^{i}}$ (19)
$+{E_{\operatorname{AC}}}\cdot(\sum\limits_{j}{\operatorname{FL}_{\operatorname{SNNConv}}^{j}\cdot f{r^{j}}+\sum\limits_{k}{\operatorname{FL}_{\operatorname{SNNFC}}^{k}\cdot f{r^{k}}}})),$
where $i$ represents every non spike-driven convolution layer and $j,k$ represent every spike-driven convolutional and fully-connected (FC) layers, respectively. Let ${E_{\operatorname{MAC}}}$ and ${E_{\operatorname{AC}}}$ denote the energy costs of MAC and AC operations, while ${FL_{\operatorname{SNNConv}}^{j}}$ and ${FL_{\operatorname{SNNFC}}^{k}}$ are the FLOPs of $j$-th convolutional layer and $k$-th FC layer. Additionally, ${f{r^{j}}}$ and ${f{r^{k}}}$ denote the firing rate (the proportion of non-zero elements) of the corresponding spike matrices at matching positions. Following the established methodology for calculating energy consumption in prior SNNs studies [38, 37, 39], we assume the data for various operations are 32-bit floating-point implementation in 45nm technology [15], in which ${E_{\operatorname{MAC}}}=4.6pJ$ and ${E_{\operatorname{AC}}}=0.9pJ$. The detailed computation methods for all metrics and the firing rate of SDTrack are presented in the Supplementary Materials.
4.3 Comparison with Standard Pipeline
Figure 4: Performance comparison of various event aggregation methods on the FE108 dataset.
Event Aggregation. We first evaluate the performance of GTP against other event aggregation methods on the FE108 dataset. As shown in Fig. 4, our method demonstrates significant advantages across five established models and SDTrack. Notably, several models that previously performed poorly in event-based tracking tasks achieve near-SOTA performance when incorporating GTP. This indicates that the inability of conventional event aggregation methods to provide sufficient temporal information is one of the key factors limiting SOTA SOT models’ effectiveness in event-based tracking tasks. This finding also validates the superior performance and generalizability of the GTP method. Furthermore, SDTrack achieves superior accuracy compared to several prominent SOT trackers, regardless of the event aggregation method employed.
Event-based Tracker. Building upon these findings, we further evaluate the effectiveness of the SDTrack architecture on event-based tracking datasets FE108 [43], VisEvent [28] and FELT [29], with detailed dataset specifications provided in the Supplementary Materials. Notably, to ensure fair comparison without the overwhelming advantages of GTP methods, most baseline approaches also utilize GTP for event aggregation, thus demonstrating SDTrack’s inherent advantages in event-based tracking tasks.
Figure 5: AUC (left) and PR (right) plot on the VisEvent dataset. Best viewed with zooming in.
As shown in Tab. 1 and Fig. 5, on all three datasets, SDTrack demonstrates SOTA potential. On the FE108 dataset, SDTrack-Tiny achieves significant improvements over the previous best model GRM, with a 4.0% increase in AUC and a 2.2% gain in PR. The Base model further enhances these performance metrics compared to its Tiny counterpart. On the FELT dataset, SDTrack outperforms STARK in AUC metrics while achieving the second-highest score in PR metrics. Notably, STARK’s PR performance relies on the GTP method; without it, STARK’s PR score drops to 50.8, comparable to SDTrack-Base. On the crucial VisEvent dataset, which presents higher complexity, SDTrack-Tiny achieves PR scores slightly below STNet and SNNTrack. However, its AUC performance remains comparable to the previous best model, SNNTrack. The larger-scale SDTrack-Base model surpasses SNNTrack’s previous state-of-the-art results, demonstrating a 2.0% improvement in AUC and a 1.1% increase in PR.
Notably, while SDTrack’s parameter count approximates that of previous Base-level models, its theoretical energy consumption on edge devices is substantially lower than models of comparable scale. Compared to HiT-B, which was specifically designed for lightweight deployment, our SDTrack-Tiny significantly outperforms across multiple datasets while utilizing less than half the parameters and energy consumption. These results substantiate both the energy efficiency characteristics of SNNs and their potential for edge deployment. Furthermore, due to the natural compatibility between SNNs and event data, coupled with the efficient design of our SDTrack pipeline, SNNs demonstrate the potential to surpass ANNs in event tracking tasks.
Figure 6: Selection of GTP hyperparameters. The entire selection process is conducted using SDTrack-Tiny on the FE108 dataset. (a) shows the selection experiment for $\alpha$, where events are aggregated into a two-channel image. (b) illustrates the selection process for $\beta$, during which the $\alpha$ in GTP is consistently set to 30.
4.4 Ablation Study
GTP. The GTP method involves two hyperparameters: $\alpha$ and $\beta$. Through experimental evaluation on the FE108 dataset, we determined that setting $\alpha$ to 30 and $\beta$ to 0.8 is the optimal configuration. The experimental process is illustrated in Fig 6.
Position Encoding. We evaluate the effectiveness of our Intrinsic Position Learning (IPL) approach. As shown in experiment 2 in Tab. 2, when IPL is not applied and the template and search images are directly fed into the SNN Conv Module (Siamese network) separately, the PR decreases by 2.04% due to lack of positional information. This demonstrates that our IPL enables the network to learn positional information effectively. Examining results from experiments 3 and 4 reveals that additional position encoding leads to performance degradation, as the network already naturally learns positional information through the IPL method, and extra positional information introduces noise. Experiments 5-7 demonstrate that the performance improvement stems from learned positional information rather than any form of joint feature extraction. These experiments conclusively demonstrate that our SDTrack does not require explicit position encoding steps, as the necessary positional information is learned inherently.
# Method AUC(%) PR(%)
1 SDTrack-Tiny 59.00 91.30
2 IPL $\rightarrow$ Siamase 58.10 89.66
3 + Learnable Positional Encoding 58.79 89.52
4 + Sinusoidal Positional Encoding 58.57 90.77
5 Intersection Size 0 $\rightarrow$ 32 58.22 90.14
6 Intersection Size 0 $\rightarrow$ 64 58.75 90.82
7 Intersection Size 0 $\rightarrow$ 128 43.91 73.34
8 Center Head $\rightarrow$ Corner Head 58.81 90.17
9 Without Pretrained 47.80 74.50
Table 2: Ablation studies on the FE108 dataset.
Tracking Head and Pre-training. Experiment 8 in Tab. 2 demonstrates that in SDTrack, the Center Head achieves superior performance compared to the Corner Head. Experiment 9 confirms that pre-training in SDTrack, consistent with other trackers, is essential.
Figure 7: Visualized comparisons of our approach and other excellent trackers OSTrack [40] and STARK [35]. Our method performs better when the target suffer from underexposure, overexposure and similar object.
5 Discussion and Limitations
Tracking in Degraded Conditions. As shown in Fig. 7, SOT trackers perform poorly under various degraded conditions, which we attribute to the lack of texture information in event data. However, SDTrack, leveraging spike signals, maintains robust target localization capability even under degraded conditions.
Compare with Other Spike-driven Backbones. As shown in Tab. 3, our SDTrack, specifically designed for tracking tasks, demonstrates significant performance advantages compared to other Spike-driven Transformer series networks. Additionally, we observe that SNN-based trackers utilizing LIF neurons struggle to adapt to tracking tasks that require precise coordinate predictions.
Backbone Neuron AUC(%) PR(%)
Spike-driven V1 [38] LIF 52.4 85.3
Spike-driven V2 [37] LIF 55.9 85.8
Spike-driven V3 [39] I-LIF 58.9 90.3
SDTrack-Tiny (Ours) I-LIF 59.0 91.3
Table 3: Comparison with candidate architectures on FE108. All settings except for the backbone are aligned with SDTrack.
Final Convolution of Tracking Head. When incorporating an additional neuronal layer before the final convolution in the tracking head, SDTrack becomes fully spike-driven. However, this modification leads to decreased performance, as shown in Tab. 4. This degradation occurs because spike signals constrain the ability of convolutional layers to output precise coordinates and target dimensions. Notably, the final convolutional layer in the tracking head operates on a small spatial scale, thus introducing only negligible additional energy consumption.
Tracking Head AUC(%) PR(%) Power(mJ)
SDTrack Head 59.0 91.3 7.9967
Spike-driven Head 58.9 90.4 7.9963
Table 4: Impact of the fully spike-driven tracking head on SDTrack performance. The experiments are conducted using SDTrack-Tiny on the FE108 dataset.
The GTP Approach Fails on ODTrack. When the GTP method is applied to SDTrack, the AUC score decreases significantly from 43.2 to 19.7. This performance degradation occurs because ODTrack learns contextual relationships using a single token, which conflicts with the inherent trajectory information encoded in the GTP method. Fortunately, we observe this issue only in this particular tracker.
6 Conclusion
In this work, we propose the first transformer-based spike-driven tracking pipeline for event-based tracking. The pipeline consists of GTP method and SDTrack. GTP method provides stronger event representation by recording trajectory information, which we demonstrate to be robust and generalizable through extensive experimentation. SDTrack is the first transformer-based spike-driven tracker that learns enhanced positional information through IPL without requiring additional positional encoding. Notably, SDTrack is end-to-end and requires no data augmentation or post-processing, yet achieves performance comparable to state-of-the-art methods across datasets while maintaining minimal energy consumption and parameter counts. We hope this work serves as a solid baseline to inspire and facilitate future research.

### v2

@@
 Third, we introduce a transformer-based spike-driven tracker, comprised of the intrinsic position learning, backbone, and head.
-3.1 I-LIF Spiking Neuron
-Recently, the emergence of I-LIF neurons enhances the expressiveness of SNNs [23, 34, 20, 25, 39].
-This enhancement enables SNNs to potentially outperform ANNs in various machie learning tasks [23, 20, 34], while preserving their spike-driven characteristics.
-Therefore, we employ the I-LIF spiking neuron model to build our SDTrack pipeline.
-I-LIF neurons transmit information through spike discharge frequency in the training stage, while utilizing binary spike signals during the inference stage [23, 39].
-In the training stage, the output of the I-LIF neuron can be expressed as
-${\mathbf{s}^{\ell}}=\frac{1}{T}\cdot\left\lfloor\operatorname{clip}\left\{{\mathbf{x}^{\ell}},\text{0},T\right\}\right\rceil,$ (1)
-[…]
-In the inference stage, the firing rate $\mathbf{s}^{\ell}$ is decoupled into $T$ binary spike signals, mathematically described as
-${\mathbf{s}^{\ell}}=\frac{1}{T}\cdot\sum\limits_{t=1}^{T}{\mathbf{s}^{\ell}[t]}.$ (2)
-At this stage, for neurons in layer $l+1$, whether considering matrix multiplication or Hadamard product as linear operators, their input can be expressed as
-${\mathbf{x}^{\ell+1}}={\mathbf{w}^{\ell}}\cdot{\mathbf{s}^{\ell}}=\sum\limits_{t=1}^{T}{\frac{{{\mathbf{w}^{\ell}}}}{T}}{\mathbf{s}^{\ell}}[t],$ (3)
-where $\mathbf{w}^{\ell}$ is the weight matrix in linear or convolutional layer, ${\mathbf{s}^{\ell}}[t]\in\{0,1\}$ is a binary spike at time $t$.
-Therefore, I-LIF maintains the key spike-driven property of SNNs during inference, which enables the transformation of MAC into efficient AC operations, significantly reducing storage and computational overhead during edge deployment.
+3.1 Unified Spiking Neuron Model
+Fundamental dynamics of postsynaptic neurons include membrane potential elevation after receiving presynaptic neuron spikes, spike generation upon membrane potential reaching threshold, and membrane potential reset following spike emission.
+Since our method is general and independent of specific neuron types, we present a unified dynamical equation for spiking neurons, described as
+$\mathbf{U}[t]$ $=\mathbf{H}[t-1]+\frac{1}{\tau}(\mathbf{X}[t]-(\mathbf{H}[t-1]-{\mathbf{U}_{reset}})),$ (1)
+$\mathbf{S}[t]$ $=f(\mathbf{U}[t]-{\mathbf{U}_{threshold}}),$ (2)
+$\mathbf{H}[t]$ $=\mathbf{U}[t](1-\mathbf{S}[t]).$ (3)
+For any neuron, $\mathbf{U}[t]$ represents its membrane potential at time $t$ after receiving input $\mathbf{X}[t]$, while $\mathbf{S}[t]$ denotes the neuron’s output at time $t$, and $\mathbf{H}[t]$ signifies the residual membrane potential following the completion of the neuron’s lifecycle at time $t$.
+The parameter $\tau$ functions as the decay factor of the neuronal membrane potential, $\mathbf{U}_{reset}$ represents the reset membrane potential post-firing, and $\mathbf{U}_{threshold}$ denotes the firing threshold of the neuron.
+These three parameters constitute the adjustable hyperparameters within the model architecture.
+Both IF (Integrate-and-Fire) [3] and LIF (Leaky Integrate-and-Fire) [23] neurons, the discrimination function $f(\cdot)$ is implemented as a Heaviside step function.
+In this formulation, when $\tau$ equals 1, the neuron functions as an IF neuron; when $\tau$ exceeds 1, it operates as a LIF neuron.
+This model represents an I-LIF neuron [37] when parameters $\mathbf{U}_{reset}$ and $\mathbf{U}_{threshold}$ are both set to 0 and $f(x)$ corresponds to ${\frac{1}{D}\cdot\operatorname{Clip}(\operatorname{round}(x),0,D)}$, where $x$ always represents $\mathbf{U}[t]$, $D$ denotes virtual time steps, $\operatorname{clip}(x,min,max)$ constrains $x$ to the range $[min,max]$, and $\operatorname{round}(\cdot)$ performs the rounding operation.
+During inference, this model can be converted to 0/1 spike signals using the spike-ahead principle, where $D$ is incorporated into actual iterative timestep $T$, forming a total of $T\times D$ time steps.
+Therefore, all these neurons achieve spike-driven implementation during the inference phase, and neurons with this characteristic replace MAC operations with simpler AC operations during inference, significantly reducing storage and computational costs for edge deployment.
 Figure 3: Example of event stream and event aggregation.
@@
 By using our GTP, the first and second channels in a event image, i.e., $h_{i}^{1}$ and $h_{i}^{2}$, accumulate the number of positive and negative polarities within the event stream, respectively.
-As shown in Fig 3, when multiple directional movements occur at a pixel within the interval $L$, prior methods only select the most recent polarity value as the aggregation result, discarding preceding motion information.
+As shown in Fig.
+3, when multiple directional movements occur at a pixel within the interval $L$, prior methods only select the most recent polarity value as the aggregation result, discarding preceding motion information.
 In contrast, GTP separately accumulates the number of positive and negative polarities, thus preserving all motion information throughout the entire time interval.
@@
 $\mathbf{U}=\operatorname{IPL}(\mathbf{X},\mathbf{Z}),$ (8)
-where $\mathbf{U}\in(T,C,{H_{z}+H_{x}},{W_{z}+W_{x}})$.
+[…]
+where $\mathbf{U}\in(T,C,{H_{z}+H_{x}},{W_{z}+W_{x}})$, ${{\rm O}_{1}}$ and ${{\rm O}_{2}}$ represent zero tensors.
 As demonstrated in Tab.
@@
 This design addresses the inherent limitation of spike-based convolutions, which are insufficient for precise object localization and scale estimation—a deficiency that cannot be compensated through continuous learning.
-We demonstrate the necessity of this approach for constructing SNN-based trackers through both experimental validation in Sec 5 and theoretical derivation in the Supplementary Materials.
-Furthermore, in Sec 4.4, we demonstrate that center heads are more suitable for SDTrack compared to corner heads.
+We demonstrate the necessity of this approach for constructing SNN-based trackers through both experimental validation in Sec.
+5 and theoretical derivation in the Supplementary Materials.
+Furthermore, in Sec.
+4.4, we demonstrate that center heads are more suitable for SDTrack compared to corner heads.
 Training and Inference.
@@
 The detailed description of the loss function is provided in the Supplementary Material.
-Since neurons transmit spike firing frequencies during the training stage, the network’s simulation time step is set to 1.
 During inference, we follow the standard SOT procedure by using the first frame of the test sequence as the template.
@@
 2.
-Notably, during inference, the network’s simulation time step is set to T, as the spike firing rate used in training is systematically decoupled into precise spike signals over T time steps according to Eq.
-2.
-Methods Parameters (M) Timestep Power (mJ) FE108 [43] FELT [29] VisEvent [28]
+Methods Param. (M) Spiking Neuron Timesteps ($T\times D$) Power (mJ) FE108 [41] FELT [28] VisEvent [27]
 AUC(%) PR(%) AUC(%) PR(%) AUC(%) PR(%)
-STARK [35] 28.23 1 58.88 57.4 89.2 39.6 51.7 34.1 46.8
-SimTrack [4] 88.64 1 93.84 56.7 88.3 36.8 47.0 34.6 47.6
-OSTrack256 [40] 92.52 1 98.90 54.6 87.1 35.9 45.5 32.7 46.4
-ARTrack256 [32] 202.56 1 174.80 56.6 88.5 39.5 49.4 33.0 43.8
-SeqTrack-B256 [6] 90.60 1 302.68 53.5 85.5 33.0 42.0 28.6 43.3
-HiT-B [17] 42.22 1 19.78 55.9 88.5 38.5 48.9 34.6 47.6
-GRM [12] 99.83 1 142.14 56.8 89.3 37.2 47.4 33.4 47.7
-HIPTrack [3] 120.41 1 307.74 50.8 81.0 38.2 48.9 32.1 45.2
-ODTrack [47] 92.83 1 335.80 43.2 69.7 29.7 35.9 24.7 34.7
-SiamRPN* [21] - 1 - - - - - 24.7 38.4
-ATOM* [9] - 1 - - - 22.3 28.4 28.6 47.4
-DiMP* [2] - 1 - - - 37.8 48.5 31.5 44.2
-PrDiMP* [10] - 1 - - - 34.9 44.5 32.2 46.9
-MixFormer* [8] 37.55 1 - - - 38.9 50.4 - -
-STNet* [44] 20.55 3 - - - - - 35.0 50.3
-SNNTrack* [46] 31.40 5 8.25 - - - - 35.4 50.4
-SDTrack-Tiny 19.61 4 8.16 59.0 91.3 39.6 50.1 35.6 49.2
-SDTrack-Base 107.26 4 30.52 59.9 91.5 39.8 50.7 37.4 51.5
+STARK [33] 28.23 - $1\times 1$ 58.88 57.4 89.2 39.3* 50.8* 34.1 46.8
+SimTrack [5] 88.64 - $1\times 1$ 93.84 56.7 88.3 36.8 47.0 34.6 47.6
+OSTrack256 [38] 92.52 - $1\times 1$ 98.90 54.6 87.1 35.9 45.5 32.7 46.4
+ARTrack256 [31] 202.56 - $1\times 1$ 174.80 56.6 88.5 39.5 49.4 33.0 43.8
+SeqTrack-B256 [7] 90.60 - $1\times 1$ 302.68 53.5 85.5 33.0 42.0 28.6 43.3
+HiT-B [18] 42.22 - $1\times 1$ 19.78 55.9 88.5 38.5 48.9 34.6 47.6
+GRM [13] 99.83 - $1\times 1$ 142.14 56.8 89.3 37.2 47.4 33.4 47.7
+HIPTrack [4] 120.41 - $1\times 1$ 307.74 50.8 81.0 38.2 48.9 32.1 45.2
+ODTrack [45] 92.83 - $1\times 1$ 335.80 43.2 69.7 29.7 35.9 24.7 34.7
+SiamRPN* [21] - - $1\times 1$ 24.7 38.4
+ATOM* [10] - - $1\times 1$ 22.3 28.4 28.6 47.4
+DiMP* [2] - - $1\times 1$ 37.8 48.5 31.5 44.2
+PrDiMP* [11] - - $1\times 1$ 34.9 44.5 32.2 46.9
+MixFormer* [9] 37.55 - $1\times 1$ 38.9 50.4 - -
+STNet* [42] 20.55 LIF $3\times 1$ 35.0 50.3
+SNNTrack* [44] 31.40 BA-LIF $5\times 1$ 8.25 - - - - 35.4 50.4
+SDTrack-Tiny 19.61 LIF $4\times 1$ 8.15 56.7 89.1 35.8 44.0 35.4 48.7
+I-LIF $2\times 2$ 9.87 55.3 88.1 35.7 45.3 35.4 49.5
+I-LIF $1\times 4$ 8.16 59.0 91.3 39.3 51.2 35.6 49.2
+SDTrack-Base 107.26 I-LIF $1\times 4$ 30.52 59.9 91.5 40.0 51.4 37.4 51.5
 Table 1: Comparison with standard SOT pipeline on three event-based tracking benchmarks.
@@
 for comparative analysis [27].
-Works marked with * indicate direct adoption of results from their respective benchmark reports.
+* denotes results directly adopted from their respective benchmark reports.
 The remaining results employ the GTP methodology for event aggregation to better demonstrate SDTrack’s performance advantages.
@@
 $E$ $=T\cdot({E_{\operatorname{MAC}}}\cdot\sum\limits_{i}{\operatorname{FL}_{\operatorname{SNNConv}}^{i}}$ (20)
-$+{E_{\operatorname{AC}}}\cdot(\sum\limits_{j}{\operatorname{FL}_{\operatorname{SNNConv}}^{j}\cdot f{r^{j}}+\sum\limits_{k}{\operatorname{FL}_{\operatorname{SNNFC}}^{k}\cdot f{r^{k}}}})),$
-where $i$ represents every non spike-driven convolution layer and $j,k$ represent every spike-driven convolutional and fully-connected (FC) layers, respectively.
-Let ${E_{\operatorname{MAC}}}$ and ${E_{\operatorname{AC}}}$ denote the energy costs of MAC and AC operations, while ${FL_{\operatorname{SNNConv}}^{j}}$ and ${FL_{\operatorname{SNNFC}}^{k}}$ are the FLOPs of $j$-th convolutional layer and $k$-th FC layer.
-Additionally, ${f{r^{j}}}$ and ${f{r^{k}}}$ denote the firing rate (the proportion of non-zero elements) of the corresponding spike matrices at matching positions.
-Following the established methodology for calculating energy consumption in prior SNNs studies [38, 37, 39], we assume the data for various operations are 32-bit floating-point implementation in 45nm technology [15], in which ${E_{\operatorname{MAC}}}=4.6pJ$ and ${E_{\operatorname{AC}}}=0.9pJ$.
+$+{E_{\operatorname{AC}}}\cdot(\sum\limits_{j}{\operatorname{FL}_{\operatorname{SNNConv\&FC}}^{j}\cdot f{r^{j}})+\sum\limits_{k}{\operatorname{FL}_{\operatorname{SSA}}^{k}}}),$
+where ${E_{\operatorname{MAC}}}$ and ${E_{\operatorname{AC}}}$ denote the energy costs of MAC and AC operations, respectively, while $\operatorname{FL}_{\operatorname{SNNConv\&FC}}$ represents the FLOPs of the $j$-th convolutional layer or FC layer, and $\operatorname{FL}_{\operatorname{SSA}}^{k}$ represents the FLOPs of the $k$-th SNN Self-Attention module.
+Additionally, ${f{r^{j}}}$ denote the firing rate (the proportion of non-zero elements) of the corresponding spike matrices at matching positions.
+Following the established methodology for calculating energy consumption in prior SNNs studies [36, 35, 37], we assume the data for various operations is implemented using 32-bit floating-point precision in 45nm technology [16], where ${E_{\operatorname{MAC}}}=4.6pJ$ and ${E_{\operatorname{AC}}}=0.9pJ$.
 The detailed computation methods for all metrics and the firing rate of SDTrack are presented in the Supplementary Materials.
@@
 Event-based Tracker.
-Building upon these findings, we further evaluate the effectiveness of the SDTrack architecture on event-based tracking datasets FE108 [43], VisEvent [28] and FELT [29], with detailed dataset specifications provided in the Supplementary Materials.
-Notably, to ensure fair comparison without the overwhelming advantages of GTP methods, most baseline approaches also utilize GTP for event aggregation, thus demonstrating SDTrack’s inherent advantages in event-based tracking tasks.
+Based on these findings, we further evaluate the effectiveness of SDTrack models of different sizes equipped with different neurons on event-based tracking datasets FE108 [41], VisEvent [27], and FELT [28], and provide detailed dataset specifications in the Supplementary Material.
+Notably, even though most baseline methods utilize GTP methods with significant performance advantages for event aggregation, SDTrack models still outperform them.
 Figure 5: AUC (left) and PR (right) plot on the VisEvent dataset. Best viewed with zooming in.
@@
 1 and Fig.
-5, on all three datasets, SDTrack demonstrates SOTA potential.
-On the FE108 dataset, SDTrack-Tiny achieves significant improvements over the previous best model GRM, with a 4.0% increase in AUC and a 2.2% gain in PR.
-The Base model further enhances these performance metrics compared to its Tiny counterpart.
-On the FELT dataset, SDTrack outperforms STARK in AUC metrics while achieving the second-highest score in PR metrics.
-Notably, STARK’s PR performance relies on the GTP method; without it, STARK’s PR score drops to 50.8, comparable to SDTrack-Base.
-On the crucial VisEvent dataset, which presents higher complexity, SDTrack-Tiny achieves PR scores slightly below STNet and SNNTrack.
-However, its AUC performance remains comparable to the previous best model, SNNTrack.
-The larger-scale SDTrack-Base model surpasses SNNTrack’s previous state-of-the-art results, demonstrating a 2.0% improvement in AUC and a 1.1% increase in PR.
-Notably, while SDTrack’s parameter count approximates that of previous Base-level models, its theoretical energy consumption on edge devices is substantially lower than models of comparable scale.
+5, across all three datasets, the SDTrack-Tiny model equipped with I-LIF neurons achieves the lowest parameter count and energy consumption.
+On the FE108 dataset, it improves AUC by 1.6% and PR by 2.0% compared to the previous best model, and demonstrates SOTA potential on the FELT and VisEvent datasets.
+The Tiny models equipped with LIF and temporally-featured I-LIF neurons ($T=2\times 2$) show SOTA potential on the FE108 and VisEvent datasets, but perform poorly on the long sequence tracking benchmark FELT dataset.
+Compared to the Tiny models, the Base models further improve these performance metrics, achieving SOTA results on both AUC and PR metrics across all three event-based tracking datasets.
+Notably, while SDTrack-Base maintains a parameter count comparable to previous Base-level models, it achieves substantially lower theoretical energy consumption on edge devices.
 Compared to HiT-B, which was specifically designed for lightweight deployment, our SDTrack-Tiny significantly outperforms across multiple datasets while utilizing less than half the parameters and energy consumption.
@@
 Figure 6: Selection of GTP hyperparameters.
-The entire selection process is conducted using SDTrack-Tiny on the FE108 dataset.
 (a) shows the selection experiment for $\alpha$, where events are aggregated into a two-channel image.
@@
 4.4 Ablation Study
+All ablation experiments follow the same experimental setup and are conducted using SDTrack-Tiny equipped with I-LIF neurons of $T=1\times 4$ on FE108.
 GTP.
@@
 Through experimental evaluation on the FE108 dataset, we determined that setting $\alpha$ to 30 and $\beta$ to 0.8 is the optimal configuration.
-The experimental process is illustrated in Fig 6.
+The experimental process is illustrated in Fig.
+6.
 Position Encoding.
@@
 9 Without Pretrained 47.80 74.50
-Table 2: Ablation studies on the FE108 dataset.
+Table 2: Ablation studies.
 Tracking Head and Pre-training.